# T2AcmrTempora-TLF

### v2

@@
Nurkhan Laiyk1  Daniil Orel1∗  Rituraj Joshi2  Maiya Goloburda1
Yuxia Wang1Preslav Nakov1Fajri Koto1
1Department of Natural Language Processing, MBZUAI
2Cerebras Systems
@@
Instruction tuning in low-resource languages remains underexplored due to limited text data, particularly in government and cultural domains. To address this, we introduce and open-source a large-scale (10,600 samples) instruction-following (IFT) dataset, covering key institutional and cultural knowledge relevant to Kazakhstan. Our dataset enhances LLMs’ understanding of procedural, legal, and structural governance topics. We employ LLM-assisted data generation, comparing open-weight and closed-weight models for dataset construction, and select GPT-4o as the backbone. Each entity of our dataset undergoes full manual verification to ensure high quality. We also show that fine-tuning Qwen, Falcon, and Gemma on our dataset leads to consistent performance improvements in both multiple-choice and generative tasks, demonstrating the potential of LLM-assisted instruction tuning for low-resource languages.
@@ 1 Introduction
Рис. 1: End-to-end process of dataset construction. The English translation is for illustrative purposes.
To demonstrate the effectiveness of this approach, we introduce an instruction-tuning dataset for Kazakh that integrates both institutional [^1] Our study incorporates administrative, procedural, legal, structural, and other government-related types of information. (GovSet) and cultural (CultSet) domains. We choose Kazakh as our case study because it remains underrepresented in NLP (Joshi et al., 2020), despite having approximately 20 million speakers. Prior research on Kazakh NLP has primarily focused on classic tasks such as named entity recognition Yeshpanov et al. (2022) and sentiment analysis Yeshpanov and Varol (2024), leaving more advanced applications like instruction tuning, largely unexplored. Our contributions are as follows:
  • We create an open-source a high-quality manually verified large scale (10K samples) IFT dataset, which covers both cultural, and institutional knowledge, relevant to Kazakhstan.
@@
  • We demonstrate that fine-tuning on our dataset, results in consistent improvements in both multiple-choice and generative tasks. These results highlight the impact of incorporating localized knowledge into instruction tuning and demonstrate the potential of LLM-assisted approaches for expanding instruction datasets in other low-resource languages.
@@ 2.1 Instruction-Tuning Datasets in English
Prior work on instruction dataset creation has largely relied on generating data from existing language models without incorporating external real-world knowledge. Self-Instruct Wang et al. (2023) expands an initial set of human-written instructions by iteratively generating new tasks using the model’s own outputs, while Honovich et al. (2023) created instruction-tuning datasets by conditioning on a few example instructions. These methods rely solely on sampling data from language models or predefined topics, rather than grounding them in external knowledge, making them less suitable for capturing domain-specific or culturally relevant understanding. Unlike these approaches, our work focuses on Kazakh, a low-resource language, and constructs an instruction dataset by leveraging external, factual sources such as governmental and cultural texts, ensuring alignment with real-world contexts.
@@ 2.2 Instruction Tuning Datasets for Medium- to Low-Resource Languages
While human-curated datasets are often expensive and require native speakers, prior work has explored automatic dataset generation using machine translation for medium- to low-resource languages. Sengupta et al. (2023) applied this approach to develop JAIS, an Arabic-centric language model, by translating various English instruction datasets into Arabic. Similarly, Li et al. (2023) translated Alpaca Taori et al. (2023) into 52 languages. More recently, Alyafeai et al. (2024) introduced an Arabic instruction dataset by translating Alpaca into Arabic and performing manual edits and localization to ensure relevance to the Arabic context. Although this method is scalable, the quality of translation remains inconsistent for low-resource languages, as noted by Li et al. (2023). Additionally, machine-translated datasets often introduce Anglocentric biases, limiting their ability to capture culturally diverse perspectives.
Several frameworks have been proposed to improve instruction tuning for low-resource languages. MURI Köksal et al. (2024) generates multilingual instruction datasets using reverse instruction generation and translation, but none of its data have been validated by native speakers. Li et al. (2024) improve upon translation-based methods by prompting LLMs in English while requiring responses in low-resource languages (e.g., Urdu), allowing models to leverage their internal knowledge of the target language’s local context. However, our approach differs in that we do not rely on the model’s internal knowledge but instead use factual information from reliable sources such as government data. Additionally, our dataset undergoes full human validation to ensure accuracy and relevance. Meanwhile, Cahyawijaya et al. (2023) focused on the linguistic aspects of instruction tuning by denoising low-resource language text and prompting models to reconstruct complete sentences. While this enhances fluency, it differs from our approach, which prioritizes grounding instructions in externally verified, domain-specific knowledge rather than refining linguistic quality alone.
@@ 3 Background
Kazakh Cultural Heritage. Kazakhstan has a rich cultural heritage that reflects a blend of nomadic traditions, Soviet influences, and modern developments.
The country’s nomadic heritage is evident in many aspects of daily life, from its architecture, with yurts (traditional felt tents) still used in rural areas, to its customs of communal gatherings and feasts, such as the celebration of the Turkic New Year, Nauryz. Kazakh hospitality is a source of pride for the nation, embodied in proverbs like "Qonaq keldi — ırysın ala keldi" ("A guest arrives, bringing fortune with them"). It is deeply important that guests are never left hungry. They are always offered a variety of traditional dishes, such as beshbarmak (a hearty dish of boiled meat with pasta) and baursak (deep-fried dough). Kazakhstan’s Soviet past has also left a lasting imprint on its culture. Many cities still bear the architectural marks of Soviet planning, while the era also shaped the country’s education and scientific institutions, fostering a strong tradition in mathematics and engineering—most notably reflected in the Baikonur Cosmodrome, the world’s first and largest space launch facility. Alongside all of this, modern developments have transformed Kazakhstan. The capital, Astana, is a prime example of this shift, with its futuristic skyline and ambitious urban projects Investments in technology, renewable energy, and digital infrastructure have propelled Kazakhstan onto the global stage, while cultural revitalization efforts have fostered a renewed interest in the Kazakh language, music, and art.
At it’s core, Kazakhstan’s culture is shaped by beliefs and values, social practices, language, artistic expression, and material culture. Understanding these components is crucial for ensuring accurate and meaningful representation.
Kazakhstan’s Institutional Structure and the Public Governance. Kazakhstan operates as a presidential republic, and since gaining independence in 1991, the country has focused on modernization, particularly in governance and legal frameworks. The Constitution of Kazakhstan, adopted in 1995, provides the foundation for the country’s laws, establishing the rights and responsibilities of citizens, as well as the structure of the government. A significant shift has been the digitalization of many governmental processes, making services more accessible to the public. Kazakhstan ranks among the top 25 countries in the UN E-Government Development Index (EGDI) (Nations, n.d.). One of the most notable achievements is the development of the eGov platform, a centralized system providing citizens with streamlined access to various government services. eGov serves as a one-stop portal for activities such as registering businesses, paying taxes, applying for social benefits, and obtaining official documents like passports and ID cards. It is a comprehensive source of governmental services in Kazakhstan, simplifying processes that previously required significant time and paperwork.
These various aspects of Kazakhstan—its rich cultural heritage, nomadic traditions, modern governance, and the increasing digitalization of public services—reflect the unique characteristics and dynamics of the country. As Kazakhstan continues to modernize and integrate technological advancements into everyday life, it is crucial that these characteristics are properly represented in tools like language models.
@@ 4.1 Document Source
We automatically collected 4,400 texts from Kazakh Wikidata,333kk.wikipedia.org specifically focusing on pages related to Kazakh culture. These pages were identified based on metadata that explicitly indicated their relevance to Kazakh cultural topics. The parsed texts include various aspects of Kazakh traditions, heritage, arts, and historical practices, providing a rich source of culturally relevant content. This ensures that the dataset reflects the depth and diversity of Kazakh culture, making it suitable for instruction-tuning tasks that require a culturally grounded perspective.
@@ 4.2 LLM-assisted Data Generation
We design a prompt (see Appendix A.3) that instructs LLMs to first extract factual information from a given Kazakh document and then generate an instruction dataset based on the extracted content. Table 2 provides detailed statistics on the source documents and the resulting instruction fine-tuning (IFT) dataset using GPT-4o. Specifically, we use 4,400 Kazakh cultural Wikipedia documents and 1,376 Kazakh government data sources, generating a total of 10,600 IFT instances. Of these, 58% belong to the government public data category (GovSet), while the remaining samples are derived from Wikipedia (CultSet). The example of generated IFT data can be found in Table 15 and Table 17.
@@
Таблица 2: Overall statistics of GPT-4o generated IFT dataset.
@@ 4.4 Final Data Overview
Table 2 further highlights a notable difference between the two subsets: the average output length in CultSet is significantly longer and contains more unique tokens than GovSet. This difference stems from the nature of GovSet responses, which are strictly factual and concise, whereas CultSet responses tend to be more diverse and expressive.
@@ 5 Experiments
We conducted two experiments: multiple-choice questions (MCQ) and text generations evaluation. We will detail each of the evaluation in the following sections.
@@
For both MCQ and generation evaluations, we use three models: Gemma-2-9B (Gemma) Team et al. (2024), Qwen-2.5-7B (Qwen)Qwen et al. (2025), and Falcon-3-10B (Falcon) Team (2024). While these LLMs have multilingual capabilities, they were not specifically trained for Kazakh, allowing us to assume that our IFT data is novel to them.
@@
We performed full fine-tuning on Gemma-2-9b, Qwen-2.5-7b, and Falcon-3-10b using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e{-5}$, and a weight decay of 0.1. We scaled the gradient norms using a maximum norm clipping value of 1.0. The learning rate was kept constant throughout the fine-tuning without any warm-up or decay with a value of 1e61e{-6} for Gemma-2 and Falcon-3, and 1e51e{-5} for Qwen-2.5. The batch size used was 16, and we packed multiple documents until the maximum sequence length was 8,192 tokens. Cross-document attention is disabled by modifying attention masks so the tokens of a document only attend to the tokens from the same document in a causal way. No adjustment were made to the original tokenizer for each model.
@@
Таблица 4: Zero-shot accuracies of language models in different datasets: (1) Datasur, (2) Constitution, and (3) Human Rights and Society
@@ 5.1 Multiple-choice Question Evaluation
A dedicated open-source Kazakh NLP community [^5] https://huggingface.co/kz-transformers has collaboratively developed and crowd-sourced multiple hand-crafted benchmarks to assess the factual knowledge of LLMs in Kazakh. We use three multiple-choice question (MCQ) datasets: (1) Dastur-MC Sagyndyk et al. (2024b), which evaluates knowledge of Kazakh traditions, (2) Kazakh Constitution-MC Sagyndyk et al. (2024a), which focuses on Kazakhstan’s legal system, and (3) Kazakh Unified National Sagyndyk et al. (2024c), which assesses citizen rights, legal protections, and societal knowledge (referred to as the "Human Rights and Society"dataset). [^6] Examples of test questions are provided in Appendix L.
@@
To assess model capability, we use the LM Eval Harness Gao et al. (2024) framework in a zero-shot setting. During evaluation, the answer is selected based on the alphabetical option with the highest likelihood.
@@
5.2 Generation Evaluation
We evaluate generation performance using our test set, which consists of 500 questions from both CultSet and GovSet (excluded from fine-tuning). We compare the best models from Section 5.1 against their vanilla counterparts. In this section, "After Fine-Tuning"refers to models fine-tuned on Alpaca + Our Data, while "Vanilla"refers to the original pre-trained models.
Automatic Evaluation with ROUGE and BERTScore
To further validate the quality of generated responses, we also evaluate BERTScore Zhang et al. . We use Kaz-RoBERTa [^7] Huggingface model: kaz-roberta-conversational as the encoder model, as it is one of the few open-source Kazakh-language transformers. The BERTScore results in Table 6 align well with the ROUGE-L scores. However, since Kazakh is a low-resource language, BERTScore should be considered a reference point rather than a definitive metric, as Kaz-RoBERTa embeddings may not perfectly capture synonym relationships.
@@ 5.2 Generation Evaluation
After FT Vanilla After FT Vanilla
Gemma 24.87 15.76 25.10 16.12
Falcon 27.98 25.96 28.70 26.17
Qwen 26.63 27.64 28.42 30.27
Таблица 5: Rouge-L comparison on CultSet and GovSet.
@@
As an extension of these experiments, we generated a set of 100 conversations for both CultSet and GovSet combined, covering topics presented in Figure 2. These conversations were intentionally left unfinished using a special prompt, as detailed in Appendix A.2. Both the original and fine-tuned models were tasked with generating the most appropriate continuation for each conversation. Examples of the resulting texts are showin in Table M. To evaluate the quality of the responses, we employed an LLM-as-a-judge framework. The results, presented in Figure 4, indicate that models fine-tuned on domain-specific data produced significantly more coherent and contextually appropriate responses compared to their pre-fine-tuning counterparts. We also see that in the conversational settings there are less ties, compared to simple question answering.
Рис. 4: Conversational data preference evaluation.

### v3

# Instruction Tuning on Public Government and Cultural Data for Low-Resource Language: a Case Study in Kazakh

@@
Nurkhan Laiyk1  Daniil Orel1∗  Rituraj Joshi3  Maiya Goloburda1
Yuxia Wang1Preslav Nakov1,2Fajri Koto1
1Mohamed bin Zayed University of Artificial Intelligence
2Institute of Foundation Models
3Cerebras Systems
@@
Instruction tuning in low-resource languages remains underexplored due to limited text data, particularly in government and cultural domains. To address this, we introduce and open-source a large-scale (10,600 samples) instruction-following (IFT) dataset, [^1] https://huggingface.co/datasets/nurkhan5l/kazakh-ift covering key institutional and cultural knowledge relevant to Kazakhstan. Our dataset enhances LLMs’ understanding of procedural, legal, and structural governance topics. We employ LLM-assisted data generation, comparing open-weight and closed-weight models for dataset construction, and select GPT-4o as the backbone. Each entity of our dataset undergoes full manual verification to ensure high quality. We also show that fine-tuning Qwen, Falcon, and Gemma on our dataset leads to consistent performance improvements in both multiple-choice and generative tasks, demonstrating the potential of LLM-assisted instruction tuning for low-resource languages.
@@ 1 Introduction
Рис. 1: Overview of the end-to-end process for constructing GovSet and CultSet datasets. English translations are for illustration purposes.
To demonstrate the effectiveness of this approach, we introduce an instruction-tuning dataset for Kazakh that integrates both institutional [^2] Our study incorporates administrative, procedural, legal, structural, and other government-related types of information. (GovSet) and cultural (CultSet) domains. We choose Kazakh as our case study because it remains underrepresented in NLP (Joshi et al., 2020), despite having approximately 20 million speakers. Prior research on Kazakh NLP has primarily focused on classic tasks such as named entity recognition Yeshpanov et al. (2022) and sentiment analysis Yeshpanov and Varol (2024), leaving more advanced applications like instruction tuning, largely unexplored.
Our contributions are as follows:
  • We create an open-source, high-quality, manually verified large scale (10K samples) IFT dataset, which covers both cultural, and institutional knowledge, relevant to Kazakhstan.
@@
  • We demonstrate that fine-tuning on our dataset results in consistent improvements in both multiple-choice and generative tasks. These results highlight the impact of incorporating localized knowledge into instruction tuning and demonstrate the potential of LLM-assisted approaches for expanding instruction datasets in other low-resource languages.
@@ 2.1 Instruction-Tuning Datasets in English
Prior work on instruction dataset creation has largely relied on generating data from existing language models without incorporating external real-world knowledge. Self-Instruct Wang et al. (2023) expands an initial set of human-written instructions by iteratively generating new tasks using the model’s own outputs, while Honovich et al. (2023) create instruction-tuning datasets by conditioning on a few example instructions. These methods rely solely on sampling data from language models or predefined topics, rather than grounding them in external knowledge, making them less suitable for capturing domain-specific or culturally relevant understanding. Unlike these approaches, our work focuses on Kazakh, a low-resource language, and constructs an instruction dataset by leveraging external, factual sources such as governmental and cultural texts, ensuring alignment with real-world contexts.
@@ 2.2 Instruction Tuning Datasets for Medium- to Low-Resource Languages
While human-curated datasets are often expensive and require native speakers, prior work has explored automatic dataset generation using machine translation for low- to medium-resource languages. Sengupta et al. (2023) applied this approach to develop JAIS, an Arabic-centric language model, by translating various English instruction datasets into Arabic. Similarly, Li et al. (2023) translated Alpaca Taori et al. (2023) into 52 languages. More recently, Alyafeai et al. (2024) introduced an Arabic instruction dataset by translating Alpaca into Arabic and then performing manual edits and localization to ensure relevance to the Arabic context. Although this method is scalable, the quality of translation remains inconsistent for low-resource languages, as noted by Li et al. (2023). Additionally, machine-translated datasets often introduce Anglocentric biases, limiting their ability to capture culturally diverse perspectives.
Several frameworks have been proposed to improve instruction tuning for low-resource languages. MURI Köksal et al. (2024) generates multilingual instruction datasets using reverse instruction generation and translation, but none of its data have been validated by native speakers. Li et al. (2024) improve upon translation-based methods by prompting LLMs in English while requiring responses in low-resource languages (e.g., Urdu), allowing models to leverage their internal knowledge of the target language’s local context. However, our approach differs by grounding responses in factual information from reliable sources, such as government data, rather than relying solely on the model’s internal knowledge. Additionally, our dataset undergoes full human validation to ensure accuracy and relevance. Meanwhile, Cahyawijaya et al. (2023) focused on the linguistic aspects of instruction tuning by denoising low-resource language text and prompting models to reconstruct complete sentences. While this enhances fluency, it differs from our approach, which prioritizes grounding instructions in externally verified, domain-specific knowledge rather than refining linguistic quality alone.
@@ 3 Background
Kazakh Cultural Heritage. Kazakhstan has a rich cultural heritage that reflects a blend of nomadic traditions, Soviet influences, and modern developments. The country’s nomadic heritage is evident in many aspects of daily life, from its architecture, with yurts (traditional felt tents) still used in rural areas, to its customs of communal gatherings and feasts, such as the celebration of the Turkic New Year, Nauryz.
Kazakhstan’s Soviet past has also left a lasting imprint on its culture. Many cities still bear the architectural marks of Soviet planning, while the era also shaped the country’s education and scientific institutions, fostering a strong tradition in mathematics and engineering—most notably reflected in the Baikonur Cosmodrome, the world’s first and largest space launch facility.
Alongside all of this, modern developments have transformed Kazakhstan. The capital, Astana, is a prime example of this shift, with its futuristic skyline and ambitious urban projects. Investments in technology, renewable energy, and digital infrastructure have propelled Kazakhstan onto the global stage, while cultural revitalization efforts have fostered a renewed interest in the Kazakh language, music, and art.
At its core, Kazakhstan’s culture is shaped by beliefs and values, social practices, language, artistic expression, and material culture. Understanding these components is crucial for ensuring accurate and meaningful representation.
Kazakhstan’s Institutional Structure and Public Governance.
Kazakhstan is a presidential republic that has prioritized modernization since its independence in 1991, particularly in governance and legal systems. The 1995 Constitution established the legal foundation, defining citizens’ rights and the structure of government. A major step in this modernization has been the digitalization of public services. Kazakhstan ranks among the top 25 countries in the UN E-Government Development Index (EGDI) (Nations, n.d.), with the eGov platform serving as a centralized portal for services like business registration, tax payments, and social benefits.
These efforts reflect a broader national context shaped by Kazakhstan’s cultural heritage, nomadic traditions, and growing digital infrastructure. Platforms like eGov highlight the integration of technology into daily governance. As the country continues to modernize, it is essential that language models accurately represent these unique characteristics to support cultural understanding and global relevance.
@@ 4.1 Document Source
We automatically collected 4,400 texts from Kazakh Wikipedia,444kk.wikipedia.org specifically focusing on pages related to Kazakh culture. These pages were identified based on metadata that explicitly indicated their relevance to Kazakh cultural topics. The parsed texts include various aspects of Kazakh traditions, heritage, arts, and historical practices, providing a rich source of culturally relevant content. This ensures that the dataset reflects the depth and diversity of Kazakh culture, making it suitable for instruction-tuning tasks that require a culturally grounded perspective.
@@ 4.2 LLM-assisted Data Generation
We design a prompt (see Appendix A.3) that instructs LLMs to first extract factual information from a given Kazakh document and then generate an instruction dataset based on the extracted content. Table 2 provides detailed statistics on the source documents and the resulting instruction fine-tuning (IFT) dataset using GPT-4o. Specifically, we use 4,400 Kazakh cultural Wikipedia documents and 1,376 Kazakh government data sources, generating a total of 10,600 IFT instances. Of these, 58% belong to the government public data category (GovSet), while the remaining samples are derived from Wikipedia (CultSet). Examples of generated IFT data can be found in Table 15 and Table 17.
@@
Таблица 2: Statistics of GPT-4o generated IFT dataset.
@@ 4.4 Final Data Overview
Table 2 further highlights a notable difference between the two subsets: the average output length in CultSet is significantly longer and includes more unique tokens than GovSet. This difference stems from the nature of GovSet responses, which are strictly factual and concise, whereas CultSet responses tend to be more diverse and expressive.
@@ 5 Experiments
We conducted two experiments: multiple-choice questions (MCQ) and text generation evaluation. We will detail each evaluation in the following sections.
@@
For both MCQ and generation evaluations, we use three instruction-tuned models: Gemma-2-9b-instruct (Gemma) Team et al. (2024), Qwen-2.5-7b-instruct (Qwen) Qwen et al. (2025), and Falcon-3-10b-instruct (Falcon) Team (2024). While these LLMs offer multilingual capabilities, none were specifically trained for Kazakh, allowing us to assume that our IFT data is novel to them.
@@
We performed full fine-tuning on Gemma-2-9b-instruct (Gemma), Qwen-2.5-7b-instruct (Qwen), and Falcon-3-10b-instruct (Falcon) using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=1e{-5}$, and a weight decay of 0.1. We scaled the gradient norms using a maximum norm clipping value of 1.0. The learning rate was kept constant throughout the fine-tuning without any warm-up or decay with a value of 1e61e{-6} for Gemma and Falcon, and 1e51e{-5} for Qwen. The batch size used was 16, and we packed multiple documents until the maximum sequence length was 8,192 tokens. Cross-document attention is disabled by modifying attention masks so the tokens of a document only attend to the tokens from the same document in a causal way. No adjustments were made to the original tokenizer for each model.
@@
Таблица 4: Zero-shot accuracies of language models in different datasets: (1) Dastur, (2) Constitution, and (3) Human Rights and Society
@@ 5.1 Multiple-choice Question Evaluation
A dedicated open-source Kazakh NLP community [^6] https://huggingface.co/kz-transformers has collaboratively developed and crowd-sourced multiple hand-crafted benchmarks to assess the factual knowledge of LLMs in Kazakh. We use three multiple-choice question (MCQ) datasets: (1) Dastur-MC Sagyndyk et al. (2024b), which evaluates knowledge of Kazakh traditions, (2) Kazakh Constitution-MC Sagyndyk et al. (2024a), which focuses on Kazakhstan’s legal system, and (3) Kazakh Unified National Sagyndyk et al. (2024c), which assesses citizen’s rights, legal protections, and societal knowledge (referred to as the "Human Rights and Society"dataset). [^7] Examples of test questions are provided in Appendix L.
@@
To assess the capabilities of the model, we use the LM Eval Harness Gao et al. (2024) framework in a zero-shot setting. During evaluation, the answer is selected based on the alphabetical option with the highest likelihood.
@@
Vanilla After FT Vanilla After FT
Gemma 15.76 24.87 16.12 25.10
Falcon 25.96 27.98 26.17 28.70
Qwen 27.64 26.63 30.27 28.42
Таблица 5: ROUGE-L comparison on CultSet and GovSet before and after fine-tuning.
@@
5.2 Generation Evaluation
We evaluate generation performance using our test set, which consists of 500 questions from both CultSet and GovSet (excluded from fine-tuning). We compare the best-performing models from Section 5.1 against their vanilla counterparts. In this section, "After Fine-Tuning"refers to models fine-tuned on Alpaca + Our Data, while "Vanilla"refers to the original instruct models.
Automatic Evaluation with ROUGE and BERTScore
To further validate the quality of generated responses, we also evaluate BERTScore Zhang et al. . We use Kaz-RoBERTa Sagyndyk et al. (2025) as the encoder model, as it is one of the few open-source Kazakh-language transformers. The BERTScore results in Table 6 align well with the ROUGE-L scores. However, since Kazakh is a low-resource language, BERTScore should be considered a reference point rather than a definitive metric, as Kaz-RoBERTa embeddings may not perfectly capture synonym relationships.
@@ 5.2 Generation Evaluation
As an extension of these experiments, we generated a set of 100 conversations for both CultSet and GovSet combined, covering topics presented in Figure 2. These conversations were intentionally left unfinished using a special prompt, as detailed in Appendix A.2. Both the original and fine-tuned models were tasked with generating the most appropriate continuation for each conversation. Examples of the resulting texts are shown in Appendix M. To evaluate the quality of the responses, we employed an LLM-as-a-judge framework. The results, presented in Figure 4, indicate that models fine-tuned on domain-specific data produced significantly more coherent and contextually appropriate responses compared to their pre-fine-tuning counterparts. We also see that in the conversational settings there are less ties, compared to simple question answering.